OpenStudy (anonymous):

expand and simplify... ∑n=04n2/2

OpenStudy (anonymous):

\[\sum_{n=0}^{4} n^2/2\]

OpenStudy (zale101):

@abb0t

OpenStudy (anonymous):

Just write it out. There are only 5 terms.

OpenStudy (anonymous):

\[ \frac {0^2} 2 + \frac {1^2}2+ \frac {2^2}2+ \frac {3^2}2+ \frac {4^2}2 \]

OpenStudy (anonymous):

Hi, @Opcode

OpenStudy (anonymous):

@wio what do I do after? like how do I simplify?

OpenStudy (anonymous):

Arithmetic.

OpenStudy (anonymous):

oh just add them together? I got 15, is that correct?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

okay thanks @wio